When 15.0 mL of a 3.27×10-4 M ammonium carbonate solution is combined with 12.0 mL of a 8.05×10-4 M manganese(II) iodide solution does a precipitate form? (yes or no) For these conditions the Reaction Quotient, Q, is equal to
Answer #1

Concentration of ammonium carbonate((NH4)2CO3) in the mixture = (3.27*10^-4)*(15/27)

                               = 1.82*10^-4 M

1 mole (NH4)2CO3 = 1 mole CO3^2-

Concentration of CO3^2- = 1.82*10^-4 M

Concentration of manganese(II) iodide(MnI2) in the mixture = (8.05*10^-4)*(12/27)

                                   = 3.58*10^-4 M

1 mole MnI2 = 1 mole Mn+2

Concentration of Mn+2 = 3.58*10^-4 M

MnCO3(s) <-----> Mn+2(aq) + CO3^2-(aq)

Ksp of MnCO3 = 1.8*10^-11

Qsp of MnCO3 = [Mn+2][CO3^2-]

Qsp of MnCO3 = (3.58*10^-4)(1.82*10^-4) = 6.5*10^-8

Qsp of MnCO3 > Ksp of MnCO3. so that, Precipitation occurs